Question 8 f(x) = 3x2 + 2 Find f(x). 0 sqrt(6x) O (3x^2 + 2)^3/2 O 1/sqrt(3x^2 + 2) O 3x/sqrt(3x^2 + 2)
Answer #1

given data:

f(x)=\sqrt{3x^2+2}

this is inthe form of

f(1) = (g())

f(2) = a* (g(2) 9-1* 9 (0)

f(x)=\sqrt{3x^2+2} =(3x^2+2)^{0.5}

f'(x)=0.5*(3x^2+2)^{-0.5}*6x

f'(x)=3x*(3x^2+2)^{-0.5}

f'(x)=\frac{3x}{\sqrt{3x^2+2}}

option-D